# Preparations for Quantum Computing in Hadron Physics

### v1

@@ 3.2 Hardware
Impressive progress in quantum computing hardware, here briefly summarized, and its availability through cloud services for researchers worldwide, is causing an explosion of interest and many exploratory works have appeared and continue to appear.
@@
Since Feynman and Deutsch initial research on quantum computers [86, 87] there has been a continuous theoretical effort in precisely defining the characteristics and requirements of fault-tolerant computations[88] as well as many experimental efforts in different hardware platforms to realize those requirements [89], we here briefly mention some of these platforms: superconducting circuits, trapped ions, linear photonic circuits and Rydberg atoms.
@@
Figure 5: Left: Energy–level scheme of a quantized harmonic oscillator $H={\rm constant}+\hbar\omega(a+a^{\dagger})$ (implementable as a microscopic oscillating LCLC circuit). Microwaves which excite the $|0\rangle\leftrightarrow|1\rangle$ transition can also excite higher levels. Right: to implement a two–level “transmon” qubit, a nonanharmonicity is needed, which is achieved by swapping the linear inductor by a non-linear Josephson junction. The Hamiltonian receives nonharmonic terms such as $\hbar\omega^{\prime}(a+a^{\dagger})^{4}$, and the equispacing is broken, so that no further excitation of the oscillator is possible with the same microwave pulse.
@@
Trapped ions[97, 81, 73] is another promising technology for quantum computing and simulation. The basic building blocks are ion chains confined in optical traps. The ion’s internal states provide the qubit’s $|0\rangle$ and $|1\rangle$. Single-qubit transformations are then realized by the excitation of individual ions with laser pulses, while entanglement can be introduced by changing the collective motions of the particles in the trap. Remarkably, these technologies give access to long-range interactions and the possbility of qudit computations [98], where the basic blocks are quantum systems with dimension higher than 2. Companies such as IonQ have commercially available systems with 25 and 36 qubits which can be accessed on the cloud; or Alpine Quantum Technologies that sells an office machine with 20 qubits. In addition to good two-qubit gate fidelity (the AQT system claims to benchmark at 0.9847(3)) and decoherence time, these platforms offer interconnectivity of all qubits (which avoids SWAP operations to control remote qubits by exchanging their states to next-neighboor ones as in other architectures).
Linear photonic circuits are another promising avenue for the development of universal quantum computation. Photons provide inherently coherent and nearly noiseless qubits, allowing single-qubit operations to be implemented with high fidelity[99]. However, this same feature makes the realization of multi-qubit gates particularly challenging. In fact, instead of relying on strong nonlinear interactions—required for direct multi-qubit operations—photonic quantum processors typically employ the so-called KLM scheme[100], where effective nonlinearities are induced through measurements, rendering entangling gates non-deterministic (presenting a finite failure probability). This is hardly acceptable for calculations, such as the ones described in this review and many others which will need millions to billions of logic gates. The probability of a flawless calculation would be negligible and the cost to obtain it would exponentiate with the number of entangling gates.
@@
Neutral atom circuits rely on arrays of neutral atoms confined in magnetic or optical traps, whose valence electrons are excited to high principal quantum number states, the so-called Rydberg states [102]. These states are characterized by large dipole moments and, being very peripheral, also large mean radii, leading to strong dipole–dipole interactions and to the Rydberg blockade phenomenon, which prevents two nearby atoms from being simultaneously excited to their Rydberg states.
Single-qubit gates are typically implemented through laser-mediated Raman transitions, in close analogy with the trapped-ion quantum computers described two paragraphs above. Two-qubit gates, on the other hand, can be realized via dipole–dipole interactions when the interaction strength is small compared to the excitation strength.
In the opposite regime—dominated by Rydberg blockade—an excited atom inhibits nearby atoms from accessing their Rydberg states, effectively serving as a control qubit. A key advantage of neutral-atom systems lies in the favorable ratio between the gate execution rate, set by the excitation strength to Rydberg states ($g\sim 10^{9}\text{ s}^{-1}$), and the dephasing rate of atoms in the trap ($\lambda\sim 0.1\text{ s}^{-1}$), which in principle allows for up to $g/\lambda\sim 10^{10}$ coherent gate operations. Nonetheless, these platforms currently face important limitations in multi-qubit gate fidelities—the highest demonstrated being $\mathcal{F}<0.8$—while fault-tolerant quantum error correction requires $\mathcal{F}>0.9999$ [103].
@@
Based on our group’s work, we can state that while the number of qubits in the superconducting platforms is now starting to be satisfactory for not too demanding applications, the number of consecutive gates that can be executed without being overran by noise has to increase.
@@ 4 Lattice-Gauge Theory on a Quantum Computer
LGT is a discretized representation of Quantum Chromodynamics on a finite spacetime grid of $n$ points with spacing $a$. In comparing with other formulations of QCD, LGT’s formalism stands out because it is most often constrained to explicitly satisfy gauge invariance. For this purpose, instead of using the gauge–dependent $A^{a}_{\mu}(y)$ fields whose quanta are the gluon partons, a change of variables [106, 107, 108] to a series of link variables $U_{n\mu}$ named Wilson lines is performed
$U_{n\mu}=\mathcal{P}\exp\left\{ig\int^{a\left(n+\hat{\mu}\right)}_{an}dy_{\nu}A_{\nu}(y)\right\}=\exp\left\{iga\sum_{\alpha}T^{\alpha}A^{\alpha}_{n\mu}\right\}$ (8)
@@
$U_{n\mu}=U\left(n,n+\hat{\mu}\right)=U\left(n+\hat{\mu},n\right)^{\dagger}=U\left(n+\hat{\mu},n\right)^{-1}$ (9)
@@
The simplest discretization of the QCD action is the Wilson action,
$S_{W}=\frac{3}{g^{2}}\sum_{n}\sum_{\mu\neq\nu}\left(1-\frac{1}{3}\text{Re}\left\{U_{n\mu\nu}\right\}\right)$ (12)
@@
where $\alpha$ and $l$ stand, respectively, for spinor and “taste” indices. The $\left(\Theta_{\rho}\right)_{\alpha l}$ matrices are constructed from Dirac’s $\gamma$ ones, and $\mathcal{N}$ is a constant to ensure the correct continuum limit. The following staggered–fermion action is then natural,
@@
where $\eta_{\mu}(\rho)=\eta_{\mu}(2n+\rho)$ is equal to $1$ or 1-1 depending on the binary values stored in $\rho$, i.e., on the lattice site within each space-time point $x$. Finally, each field $\chi_{\rho}(n)$ can be promoted with a color index so under a gauge transformation $\chi_{\rho}(n)\rightarrow\Omega_{n}\chi_{\rho}(n)$, rendering gauge invariant the following lattice action,
@@ 4.1 Hamiltonian formulations
Kogut & Sussking (KS) Hamiltonian formulation of non-Abelian LGTs is based on an elegant rigid rotator analogy [110]. The Weyl-gauge ($A^{0}=0$) Hamiltonian corresponding to Eq. (15) is [110, 111]
@@
Key towards the simulation of Eq. (16) is the definition of a mapping between the Hilbert space of Chromodynamics and that of the quantum–computer as mentioned in section 3.2 above.
@@ 4.2.2 Gauge-boson mappings
The computer-discretized linear space corresponding to the gauge fields in Eq. (16) arises from direct products of single-link Hilbert spaces. We review here a construction in the electric-field basis due to Byrnes and Yamamoto[62], although several other basis are possible [59].
@@
Link states are characterized by two sets of “angular” quantum numbers, one for each of the link’s ends. For $SU(3)$ we have
@@
where $p,q$ label the irreducible representation of $SU(3)$ and $\lambda_{L/R}$ is a shorthand for the quantum numbers that label the left and right ends of the link respectively. These quantum numbers can be defined in different ways; we here employ the representation of the color group mapped from the flavor ($u$, $d$, $s$) triplet, adopting the names “isospin” $T$ and “hypercharge” $Y$. The explicit state of the link is therefore
@@
The Wilson line operators are characterized by the link quantum numbers and they are either in the fundamental ($p=1,q=0$) or adjoint ($p=0,q=1$) representations. Single-link states are therefore:
@@
Next, the scaling of the number of operations depends on the configuration of qubits in the actual device; if there were only nearest–neighbor interactions the scaling would be $\mathcal{O}(M^{2}d^{2}D)$, i.e., quadratic in the lattice–site number, while with long-range interactions it can be reduced to be linear in the lattice sites $\mathcal{O}(Md)$.
@@ 5 Particle-based encoding of Chromodynamics
The encoding reviewed in this section [116] is formulated from a few-body perspective in which the basic object is the particle. The quantum computer memory is organized in such a way that it resembles quantum field theory in the particle basis in second quantization.
@@
In contrast to particle encodings mentioned in the previous section 4.2 (among others), this encoding implements creation and annihilation operators that fulfill both commutation and anti-commutation relations exactly up to boundary terms that emerge when the memory is completely full (and provided the memory is filled following a certain order). In particular, this implies that the Pauli principle among fermions is automatically implemented, but without restricting bosons.
@@ 5.1.1 Single-particle boson operators
For a particle with momentum, spin and color quantum numbers, the vacuum state is represented by the register:
$|\Omega\rangle\equiv|0\rangle_{P/A}\otimes|0\rangle_{\text{spin}}\otimes|00\rangle_{\text{color}}\otimes|0\dots 0\rangle_{\text{momentum}}$ (26)
Encoding of spin and color requires 1 and 2 qubits, respectively, whereas the $N_{p}$ values of 1-dimensional discretized momentum require $\log_{2}N_{p}$ qubits $(\log_{2}N_{p})^{3}$ in three dimensions].
@@
The control operators therein act on the presence/absence qubit and fulfill the following relations
@@
When acting on the vacuum state $|\Omega\rangle$, Eq. (27) yields a one-particle state with spin $s$, color $c$ and momentum $p$:
$a^{\dagger}_{s,c,p}|\Omega\rangle$ $\equiv$ $\mathfrak{C}_{10}|0\rangle_{P/A}\otimes\mathfrak{s}_{s}^{\dagger}|0\rangle_{s}\otimes\mathfrak{s}_{c}^{\dagger}|0\rangle_{c}\otimes\mathfrak{s}_{p}^{\dagger}|0\dots 0\rangle_{p}$ (29)
$=$ $|1\rangle_{P/A}\otimes|s\rangle\otimes|c\rangle\otimes|p\rangle\ =\ |1scp\rangle\ .$
@@ 5.1.2 Two-particle boson operators
$|\Omega\rangle\ =\ |\Omega\rangle_{2}\otimes|\Omega\rangle_{1}\equiv|0\rangle_{P/A}|0\dots 0\rangle_{\text{momentum}}\otimes|0\rangle_{P/A}|0\dots 0\rangle_{\text{momentum}}$ (30)
where spin and color quantum numbers are now hidden for simplicity of reading.
@@
$a_{q_{1}}^{(2)\dagger}|\Omega\rangle=|q_{1}\rangle\equiv|\Omega\rangle_{2}\otimes(|1\rangle|q_{1}\rangle)_{1}=|\Omega\rangle_{2}\otimes|1q_{1}\rangle_{1}\ .$ (31)
@@
As a consequence of the limited computer memory (there, only two registers), the Fock space is truncated, introducing a discretization error in the second quantized problem. With an increasing number of particle registers, this inconvenience is pushed up far from the few–body spectrum of the theory.
@@
The creation operator is then a sum of the first suboperator plus the second one with a symmetrizer (since it produces a doubly-occupied memory),
@@
$\frac{1}{\sqrt{2}}\left({\text{$\mathbb{I}$}}\otimes{\text{$\mathbb{I}$}}+\mathcal{P}_{21}\right)$ $\equiv$ $\mathcal{S}_{2}$ (37)
a 2-particle symmetrizer. The operators then verfy the boson commutation relations
@@
when acting on the vacuum and one particle states, but they fail when the memory is full, as trying to create a third particle on the two-particle register is not possible by construction, so a creation operator acting on two-particle operators yields 0 as per Eq. (33). Therefore, the correct commutation relations are fulfilled up to a boundary term that activates when the memory is totally filled, in this case by two bosons.

### v2

@@ 3.2 Hardware
Impressive progress in quantum computing hardware, here briefly summarised, and its availability through cloud services for researchers worldwide, is causing an explosion of interest and many exploratory works have appeared and continue to appear.
@@
Since Feynman and Deutsch initial research on quantum computers [86, 87] there has been a continuous theoretical effort in precisely defining the characteristics and requirements of fault-tolerant computations[88] as well as many experimental efforts in different hardware platforms to realise those requirements [89], we here briefly mention some of these platforms: superconducting circuits, trapped ions, linear photonic circuits and Rydberg atoms.
@@
Figure 5: Left: Energy–level scheme of a quantised harmonic oscillator $H={\rm constant}+\hbar\omega(a+a^{\dagger})$ (implementable as a microscopic oscillating LCLC circuit). Microwaves which excite the $|0\rangle\leftrightarrow|1\rangle$ transition can also excite higher levels. Right: to implement a two–level “transmon” qubit, a nonanharmonicity is needed, which is achieved by swapping the linear inductor by a non-linear Josephson junction. The Hamiltonian receives nonharmonic terms such as $\hbar\omega^{\prime}(a+a^{\dagger})^{4}$, and the equispacing is broken, so that no further excitation of the oscillator is possible with the same microwave pulse.
@@
Trapped ions[97, 81, 73] is another promising technology for quantum computing and simulation. The basic building blocks are ion chains confined in optical traps. The ion’s internal states provide the qubit’s $|0\rangle$ and $|1\rangle$. Single-qubit transformations are then realised by the excitation of individual ions with laser pulses, while entanglement can be introduced by changing the collective motions of the particles in the trap. Remarkably, these technologies give access to long-range interactions and the possibility of qudit computations [98], where the basic blocks are quantum systems with dimension higher than 2. Companies such as IonQ have commercially available systems with 25 and 36 qubits which can be accessed on the cloud; or Alpine Quantum Technologies that sells an office machine with 20 qubits. In addition to good two-qubit gate fidelity (the AQT system claims to benchmark at 0.9847(3)) and decoherence time, these platforms offer interconnectivity of all qubits (which avoids SWAP operations to control remote qubits by exchanging their states to next-neighbour ones as in other architectures).
Linear photonic circuits are another promising avenue for the development of universal quantum computation. Photons provide inherently coherent and nearly noiseless qubits, allowing single-qubit operations to be implemented with high fidelity[99]. However, this same feature makes the realisation of multi-qubit gates particularly challenging. In fact, instead of relying on strong nonlinear interactions—required for direct multi-qubit operations—photonic quantum processors typically employ the so-called KLM scheme[100], where effective nonlinearities are induced through measurements, rendering entangling gates non-deterministic (presenting a finite failure probability). This is hardly acceptable for calculations, such as the ones described in this review and many others which will need millions to billions of logic gates. The probability of a flawless calculation would be negligible and the cost to obtain it would exponentiate with the number of entangling gates.
@@
Neutral atom circuits rely on arrays of neutral atoms confined in magnetic or optical traps, whose valence electrons are excited to high principal quantum number states, the so-called Rydberg states [102]. These states are characterised by large dipole moments and, being very peripheral, also large mean radii, leading to strong dipole–dipole interactions and to the Rydberg blockade phenomenon, which prevents two nearby atoms from being simultaneously excited to their Rydberg states.
Single-qubit gates are typically implemented through laser-mediated Raman transitions, in close analogy with the trapped-ion quantum computers described two paragraphs above. Two-qubit gates, on the other hand, can be realised via dipole–dipole interactions when the interaction strength is small compared to the excitation strength.
In the opposite regime—dominated by Rydberg blockade—an excited atom inhibits nearby atoms from accessing their Rydberg states, effectively serving as a control qubit. A key advantage of neutral-atom systems lies in the favourable ratio between the gate execution rate, set by the excitation strength to Rydberg states ($g\sim 10^{9}\text{ s}^{-1}$), and the dephasing rate of atoms in the trap ($\lambda\sim 0.1\text{ s}^{-1}$), which in principle allows for up to $g/\lambda\sim 10^{10}$ coherent gate operations. Nonetheless, these platforms currently face important limitations in multi-qubit gate fidelities—the highest demonstrated being $\mathcal{F}<0.8$—while fault-tolerant quantum error correction requires $\mathcal{F}>0.9999$ [103].
@@
Based on our group’s work, we can state that while the number of qubits in the superconducting platforms is now starting to be satisfactory for not too demanding applications, the number of consecutive gates that can be executed without being overrun by noise has to increase.
@@
Because implementing these many degrees of freedom is very expensive, our and other’s calculations are very simple, limited to a few particles, and few momenta. The focus has been on testing and developing algorithms for a future quantum computer, not on the classical simulations per se.
@@ 4 Lattice-Gauge Theory on a Quantum Computer
LGT is a discretised representation of Quantum Chromodynamics on a finite spacetime grid of $n$ points with spacing $a$. In comparing with other formulations of QCD, LGT’s formalism stands out because it is most often constrained to explicitly satisfy gauge invariance. For this purpose, instead of using the gauge–dependent $A^{a}_{\mu}(y)$ fields whose quanta are the gluon partons, a change of variables [106, 107, 108] to a series of link variables $U_{n\mu}$ named Wilson lines is performed
$U_{n\mu}=\mathcal{P}\exp\left\{ig\int^{a\left(n+\hat{\mu}\right)}_{an}dy_{\nu}A_{\nu}(y)\right\}=\exp\left\{iga\sum_{\alpha}T^{\alpha}A^{\alpha}_{n\mu}\right\},$ (8)
@@
$U_{n\mu}=U\left(n,n+\hat{\mu}\right)=U\left(n+\hat{\mu},n\right)^{\dagger}=U\left(n+\hat{\mu},n\right)^{-1},$ (9)
@@
The simplest discretisation of the QCD action is the Wilson action
$S_{W}=\frac{3}{g^{2}}\sum_{n}\sum_{\mu\neq\nu}\left(1-\frac{1}{3}\text{Re}\left\{U_{n\mu\nu}\right\}\right),$ (12)
@@
where $\alpha$ and $l$ stand, respectively, for spinor and “taste” indices. The $\left(\Theta_{\rho}\right)_{\alpha l}$ matrices are constructed from Dirac’s $\gamma$, and $\mathcal{N}$ is a constant to ensure the correct continuum limit. The following staggered–fermion action is then natural,
@@
where $\eta_{\mu}(\rho)=\eta_{\mu}(2n+\rho)$ is equal to $1$ or 1-1 depending on the binary values stored in $\rho$, i.e., on the lattice site within each space-time point $x$. Finally, each field $\chi_{\rho}(n)$ can be promoted with a colour index so under a gauge transformation $\chi_{\rho}(n)\rightarrow\Omega_{n}\chi_{\rho}(n)$, rendering gauge invariant the following lattice action,
@@ 4.1 Hamiltonian formulations
Kogut & Sussking (KS) Hamiltonian formulation of non-Abelian LGTs (HLGT) is based on an elegant rigid rotator analogy [110]. The Weyl-gauge ($A^{0}=0$) Hamiltonian corresponding to Eq. (15) is [110, 111]
@@
Key towards the simulation of Eq. (16) is the definition of a mapping between the Hilbert space of Chromodynamics and that of the quantum–computer as described in the next subsection.
@@ 4.2.2 Gauge-boson mappings
The computer-discretised linear space corresponding to the gauge fields in Eq. (16) arises from direct products of single-link Hilbert spaces. We review here a construction in the electric-field basis due to Byrnes and Yamamoto[62], although several other basis are possible [59].
@@
Link states are characterised by two sets of “angular” quantum numbers, one for each of the link’s ends. For $SU(3)$ we have
@@
where $p,q$ label the irreducible representation of $SU(3)$ and $\lambda_{L/R}$ is a shorthand for the quantum numbers that label the left and right ends of the link respectively. These quantum numbers can be defined in different ways; we here employ the representation of the colour group mapped from the flavour ($u$, $d$, $s$) triplet, adopting the names “isospin” $T$ and “hypercharge” $Y$. The explicit state of the link is therefore
@@
The Wilson line operators are characterised by the link quantum numbers and they are either in the fundamental ($p=1,q=0$) or adjoint ($p=0,q=1$) representations. Single-link states are therefore:
@@
Next, the scaling of the number of operations depends on the configuration of qubits in the actual device; if there were only nearest–neighbour interactions the scaling would be $\mathcal{O}(M^{2}d^{2}D)$, i.e., quadratic in the lattice–site number, while with long-range interactions it can be reduced to be linear in the lattice sites $\mathcal{O}(Md)$.
@@ 5 Particle-based encoding of Chromodynamics
The encoding reviewed in this section [116] is formulated from a few-body perspective in which the basic object is the particle. The quantum computer memory is organised in such a way that it resembles quantum field theory in the particle basis in second quantisation.
@@
In contrast to particle encodings mentioned in the previous section 4.2 (among others), this encoding implements creation and annihilation operators that fulfil both commutation and anti-commutation relations up to boundary terms that emerge when the memory is completely occupied (and provided the memory is filled following a certain order). In particular, this implies that the Pauli principle among fermions is automatically implemented, but without restricting bosons.
@@ 5.1.1 Single-particle boson operators
For a particle with momentum, spin and colour quantum numbers, the vacuum state is represented by the register:
$|\Omega\rangle\equiv|0\rangle_{P/A}\otimes|0\rangle_{\text{spin}}\otimes|00\rangle_{\text{colour}}\otimes|0\dots 0\rangle_{\text{momentum}}$ (26)
Encoding of spin and colour requires 1 and 2 qubits, respectively, whereas the $N_{p}$ values of 1-dimensional discretised momentum require $\log_{2}N_{p}$ qubits $(\log_{2}N_{p})^{3}$ in three dimensions].
@@
The control operators therein act on the presence/absence qubit and fulfil the following relations
@@
When acting on the vacuum state $|\Omega\rangle$, Eq. (27) yields a one-particle state with spin $s$, colour $c$ and momentum $p$:
$a^{\dagger}_{s,c,p}|\Omega\rangle$ $\equiv$ $|1\rangle_{P/A}\otimes|s\rangle\otimes|c\rangle\otimes|p\rangle\ =\ |1scp\rangle\ .$ (29)
@@ 5.1.2 Two-particle boson operators
$|\Omega\rangle\ =\ |\Omega\rangle_{2}\otimes|\Omega\rangle_{1}\equiv|0\rangle_{P/A}|0\dots 0\rangle_{\text{momentum}}\otimes|0\rangle_{P/A}|0\dots 0\rangle_{\text{momentum}}\ ,$ (30)
where spin and colour quantum numbers are now hidden for simplicity of reading.
@@
$a_{q_{1}}^{(2)\dagger}|\Omega\rangle=|q_{1}\rangle\equiv|\Omega\rangle_{2}\otimes|1q_{1}\rangle_{1}\ .$ (31)
@@
As a consequence of the limited computer memory (only two registers), the Fock space is truncated, introducing a discretisation error in the second quantised problem. With an increasing number of particle registers, this inconvenience is pushed up far from the few–body spectrum of the theory.
@@
The creation operator is then a sum of the first suboperator plus the second one with a symmetriser (since it produces a doubly-occupied memory),
@@
$\frac{1}{\sqrt{2}}\left({\text{$\mathbb{I}$}}\otimes{\text{$\mathbb{I}$}}+\mathcal{P}_{21}\right)$ $\equiv$ $\mathcal{S}_{2}\ ,$ (37)
a 2-particle symmetriser. The operators then verfy the boson commutation relations
@@
when acting on the vacuum and one particle states, but they fail when the memory is full, as trying to create a third particle on the two-particle register is not possible by construction, so a creation operator acting on two-particle operators yields 0 as per Eq. (33). Therefore, the correct commutation relations are fulfiled up to a boundary term that activates when the memory is totally filled, in this case by two bosons.